The perimeter of a triangle is 195 mm. If the lengths of the sides are consecutive odd integers, find the length of each side.

Answer by vleith(2983)

Perimeter = a+b+c
b=a+2, c=a+4
Perimeter = 195 = a + a+2 + a+4 = 3a + 6
3a = 189
a = 63
b = 65
c = 67